Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Board of directors

Expert replies
by chieftang » Sun Dec 04, 2011 7:52 pm
Thought I'd throw this one up for you guys. Adapted from an OG question:

A board of directors has 10 members, including Gonzalo. Randomly, one of the 10 members is chosen to be the president, one of the remaining 9 members is randomly chosen to be the treasurer, and one of the remaining 8 members is randomly chosen to be the secretary. What is the probability that Gonzalo will be either the member chosen to be the secretary or the member chosen to be the treasurer?

(A) 1/720
(B) 1/20
(C) 1/10
(D) 1/9
(E) 1/5

OA to come.
Join the discussion
Source: — Problem Solving |

by user123321 » Sun Dec 04, 2011 9:45 pm
[spoiler]IMO 1/5[/spoiler]

(9*1*8)/(10*9*8) + (9*8*1)/(10*9*8)

user123321
Just started my preparation :D
Want to do it right the first time.
Join the discussion

by GMATGuruNY » Mon Dec 05, 2011 4:57 am
chieftang wrote:Thought I'd throw this one up for you guys. Adapted from an OG question:

A board of directors has 10 members, including Gonzalo. Randomly, one of the 10 members is chosen to be the president, one of the remaining 9 members is randomly chosen to be the treasurer, and one of the remaining 8 members is randomly chosen to be the secretary. What is the probability that Gonzalo will be either the member chosen to be the secretary or the member chosen to be the treasurer?

(A) 1/720
(B) 1/20
(C) 1/10
(D) 1/9
(E) 1/5

OA to come.
P(A or B) = P(A) + P(B).

Each of the 10 members has a 1/10 chance of being chosen secretary or treasury.
The intervening elections are IRRELEVANT.

P(Gonzales is chosen secretary or treasurer) = 1/10 + 1/10 = 1/5.

The correct answer is E.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by chieftang » Mon Dec 05, 2011 7:45 am
GMATGuruNY wrote:
chieftang wrote:Thought I'd throw this one up for you guys. Adapted from an OG question:

A board of directors has 10 members, including Gonzalo. Randomly, one of the 10 members is chosen to be the president, one of the remaining 9 members is randomly chosen to be the treasurer, and one of the remaining 8 members is randomly chosen to be the secretary. What is the probability that Gonzalo will be either the member chosen to be the secretary or the member chosen to be the treasurer?

(A) 1/720
(B) 1/20
(C) 1/10
(D) 1/9
(E) 1/5

OA to come.
P(A or B) = P(A) + P(B).

Each of the 10 members has a 1/10 chance of being chosen secretary or treasury.
The intervening elections are IRRELEVANT.

P(Gonzales is chosen secretary or treasurer) = 1/10 + 1/10 = 1/5.

The correct answer is E.
Yes!!

And it bears out when you multiply the odds of NOT being chosen as president by the odds of being chosen as treasurer: (9/10)(1/9) = 1/10, and the odds of not being chosen as president, not being chosen as treasurer, and being chosen as secretary: (9/10)(8/9)(1/8) = 1/10

OA [spoiler]E[/spoiler]
Join the discussion

by gunjan1208 » Fri Dec 09, 2011 5:43 am
Hi,
I did it exactly the way described by Cheftang and got 1/10...

But the language does say that remaining 9...Remaining 8 etc....Therefore it needs to be done that way. What am I missing?

Guys, help.
Join the discussion

by rijul007 » Sat Dec 10, 2011 11:40 am
chieftang wrote:Thought I'd throw this one up for you guys. Adapted from an OG question:

A board of directors has 10 members, including Gonzalo. Randomly, one of the 10 members is chosen to be the president, one of the remaining 9 members is randomly chosen to be the treasurer, and one of the remaining 8 members is randomly chosen to be the secretary. What is the probability that Gonzalo will be either the member chosen to be the secretary or the member chosen to be the treasurer?

(A) 1/720
(B) 1/20
(C) 1/10
(D) 1/9
(E) 1/5

OA to come.
Total no of ways = 10*9*8
No of possible selections
i) when Gonzalo is the treasurer = 9*1*8
ii) when he is the secr = 9*8*1
Total no of selections = 2*9*8

Probability = 2*9*8/10*9*8 = 1/5

Option E
Join the discussion